# A central limit theorem for a generalization of the Ewens measure to random tuples of commuting permutations

### v2

@@ 1.1. The main result
For $n\geq 0$, the symmetric group $\mathfrak{S}_{n}$ comes with a natural left action $(\sigma,i)\mapsto\sigma(i)$ on the set $n]:=\{1,2,\ldots,n\}$. If $H$ is a subgroup of $\mathfrak{S}_{n}$, then the restriction of this action to $H$ gives us access to the enumerative quantity $\kappa_{n}(H):=|[n]/H|$, where we used $|\cdot|$ to denote the cardinality of finite sets, and where $[n]/H$ is the set of orbits for the action of $H$ on $[n]$. We will be particularly interested in the situation where $H=\langle\sigma_{1},\ldots,\sigma_{\ell}\rangle$ is the subgroup generated by a random ordered $\ell$-tuple $(\sigma_{1},\ldots,\sigma_{\ell})$ of pairwise commuting permutations, for some fixed $\ell\geq 1$. Our main result is a central limit theorem for the number of joint orbits $\kappa_{n}(\langle\sigma_{1},\ldots,\sigma_{\ell}\rangle)$ or simply $\kappa_{\ell,n}(\sigma_{1},\ldots,\sigma_{\ell})$. More precisely, let us introduce the set of commuting tuples of permutations
@@
This is the PMF of the much studied Ewens distribution on the symmetric group, which was introduced in the applied context of mathematical biology [18], but is also an important notion from a pure mathematics point of view (see [35, 14, 40] for insightful reviews, and the book [6] for an in-depth introduction to the Ewens measure and many other related topics in probability theory and combinatorics). The uniform $x=1$ case was of course studied much earlier than the article by Ewens, but we also note that the particular weight $x=2$ per cycle (in a more difficult spatially dependent rather than mean field situation) features in the work of Tóth on quantum spin systems [42, §5].
@@ 2. Preliminaries on the $Z$ functions
In order to obtain the leading asymptotics for the $Z_{\alpha_{1},\ldots,\alpha_{\ell}}^{[\ell]}(t)$ when $t\rightarrow 0^{+}$, we need the following elementary lemma which controls the Euler-Maclaurin approximation of integrals on $(0,\infty)$ which are possibly improper near zero.
@@
Let $f:(0,\infty)\rightarrow\mathbb{C}$ be a differentiable function. Suppose there exist $\gamma,c_{1},c_{2}>0$, such that, for all $t>0$,
$\max(|f(t)|,t|f^{\prime}(t)|)\leq c_{1}t^{\gamma-1}e^{-c_{2}t}\ .$
Then $f$ is integrable on $(0,\infty)$, the series $\sum_{\delta=1}^{\infty}f(\delta t)$ converges absolutely for all $t>0$, and we have
@@
Proof: The integrability and the convergence of the series are obvious from the given bound on $|f(t)|$. Define
$\Delta(t)$ $:=$ $\int_{t}^{\infty}f(u)\ {\rm d}u-t\sum_{\delta=1}^{\infty}f(\delta t)$
$=$ $\sum_{\delta=1}^{\infty}\int_{(\delta t,(\delta+1)t]}[f(u)-f(\delta t)]\ {\rm d}u\ .$
By the mean value theorem,
$f(u)-f(\delta t)=(u-\delta t)\times f^{\prime}(\xi)\ ,$
for some $\xi$ with $\delta t<\xi<u\leq(\delta+1)t$. For such a $\xi$, we must have
$|f^{\prime}(\xi)|\leq c_{1}\xi^{\gamma-2}e^{-c_{2}\xi}\leq c_{1}\xi^{\gamma-2}e^{-c_{2}\delta t}\ ,$
by the bound on $f^{\prime}$ given as hypothesis. Since we do not know the sign of $\gamma-2$, i.e., do not know if the power law with that exponent is increasing or decreasing, we just write
$\xi^{\gamma-2}\leq\max\left((\delta t)^{\gamma-2},\left((\delta+1)t\right)^{\gamma-2}\right)=(\delta t)^{\gamma-2}\max\left(1,\left(1+\frac{1}{\delta}\right)^{\gamma-2}\right)\ .$
If $\gamma-2\geq 0$, then $\left(1+\frac{1}{\delta}\right)^{\gamma-2}\leq 2^{\gamma-2}$, because $\delta\geq 1$. If $\gamma-2<0$, then $\left(1+\frac{1}{\delta}\right)^{\gamma-2}\leq 1$. In all cases we have
$\xi^{\gamma-2}\leq\max(1,2^{\gamma-2})\times(\delta t)^{\gamma-2}\ .$
We put absolute values/moduli and use these estimates to bound $\Delta(t)$ by
$|\Delta(t)|\leq\sum_{\delta=1}^{\infty}c_{1}\times\max(1,2^{\gamma-2})\times(\delta t)^{\gamma-2}\times e^{-c_{2}\delta t}\times\int_{\delta t}^{(\delta+1)t}(u-\delta t)\ {\rm d}u\ .$
The remaining integrals all being equal to $\frac{t^{2}}{2}$, we obtain
$|\Delta(t)|\leq\frac{c_{1}}{2}\max(1,2^{\gamma-2})t^{\gamma}\sum_{\delta=1}^{\infty}\delta^{\gamma-2}e^{-c_{2}\delta t}=\frac{c_{1}}{2}\max(1,2^{\gamma-2})t^{\gamma}\ Z_{\gamma-1}^{[1]}(c_{2}t)\ .$
Since $\gamma>0$ by hypothesis, we have $\gamma>\max(0,\gamma-1)$, and it is possible to pick some $\beta$ such that $\gamma>\beta>\max(0,\gamma-1)$. Using Lemma 2.1, while tossing the exponential decay, we can bound $Z_{\gamma-1}^{[1]}(c_{2}t)$ by a constant times $t^{-\beta}$. As a result, $|\Delta(t)|$ is bounded by a constant times $t^{\gamma-\beta}$ which goes to zero as $t\rightarrow 0^{+}$. Since $f$ is integrable on $(0,\infty)$, the integral $\int_{0}^{t}f(u)\ {\rm d}u$ also goes to zero as $t\rightarrow 0^{+}$, and the lemma follows. ∎
@@
The functions $f$ is differentiable, and we will show that it satisfies the required bounds for the application of Lemma 2.2. By hypothesis, ${\rm Re}\ \alpha_{1}>\max(0,{\rm Re}\ \alpha_{2})$, and it is therefore possible to choose some real number $\beta$ such that ${\rm Re}\ \alpha_{1}>\beta>\max(0,{\rm Re}\ \alpha_{2})$. We now use Lemma 2.1 which provides constants $c_{1,1},c_{1,2}>0$ such that, for all $t>0$,
@@
Since our chosen $\beta$ also satisfies $\beta+1>\max(0,{\rm Re}(\alpha_{2})+1)$, Lemma 2.1 again provides us with constants $c_{2,1},c_{2,2}>0$ such that for all $t>0$,
$\left|Z_{\alpha_{2}+1,\ldots,\alpha_{\ell}+1}^{[\ell-1]}(t)\right|\leq c_{2,1}t^{-\beta-1}e^{-c_{2,2}t}\ .$
Note that by (9) we have
$f^{\prime}(t)=(\alpha_{1}-1)t^{\alpha_{1}-2}Z_{\alpha_{2},\ldots,\alpha_{\ell}}^{[\ell-1]}(t)-t^{\alpha_{1}-1}Z_{\alpha_{2}+1,\ldots,\alpha_{\ell}+1}^{[\ell-1]}(t)\ .$
We then easily see that the estimates needed for the use of Lemma 2.2 hold with $\gamma:={\rm Re}(\alpha_{1})-\beta>0$ and the constants
$c_{1}$ $=$ $\max(c_{1,1},|\alpha_{1}-1|c_{1,1}+c_{2,1})\ ,$
$c_{2}$ $=$ $\min(c_{1,2},c_{2,2})\ .$
Hence, Lemma 2.2 shows that
@@ 3. Preparation for the saddle point analysis
Now that (5) is justified with $r=e^{-t}$ for some $t>0$ to be chosen shortly, we look at the modulus of the integrand $|z^{-n}\mathcal{G}_{\ell}(x,z)|=\exp(nt+x{\rm Re}\ \mathcal{L}_{\ell}(z))$. For $z=e^{-t+i\theta}$, with $\pi<\theta<\pi$, we have
@@ 4.3. Putting the estimates together and deriving the $\mathcal{J}$ integral asymptotics
involving the sequence of functions $f_{n}:\mathbb{R}\rightarrow\mathbb{C}$ define by
@@
$\lim\limits_{n\rightarrow\infty}\int_{\mathbb{R}}f_{n}(\Theta)\ {\rm d}\Theta=\int_{\mathbb{R}}e^{is\sqrt{\ell-1}-\frac{\Theta^{2}}{2}}\ {\rm d}\Theta=\sqrt{2\pi}\ e^{-\frac{(\ell-1)s^{2}}{2}}\ ,$
@@ 5. Completion of the proofs of the main proposition and theorem
Acknowledgements: A. A. thanks Bernhard Heim and Markus Neuhauser for instructive discussions related to the log-concavity of $\mathcal{H}_{\ell,n}(1)$ and related Bessenrodt-Ono inequalities. A. A. also thanks Ofir Gorodetsky and Mark Wildon for enlighting comments or posts on MathOverflow related to the asymptotics of $Z_{1,0}^{[2]}(t)$. We again thank Ofir Gorodetsky for explaining to us the connection of our work to the CLT by Maples, Nikeghbali and Zeindler and the mean asymptotics by Ercolani and Ueltschi. S. S. gratefully acknowledges a Simons collaboration grant. S. S. also thanks Jim Pitman for pointing out useful references.

### v3

@@ 1.1. The main result
For $n\geq 0$, the symmetric group $\mathfrak{S}_{n}$ comes with a natural left action $(\sigma,i)\mapsto\sigma(i)$ on the set $n]:=\{1,2,\ldots,n\}$. If $H$ is a subgroup of $\mathfrak{S}_{n}$, then the restriction of this action to $H$ gives us access to the enumerative quantity $\kappa_{n}(H):=|[n]/H|$, where we used $|\cdot|$ to denote the cardinality of finite sets, and where $[n]/H$ is the set of orbits for the action of $H$ on $[n]$. We will be particularly interested in the situation where $H=\langle\sigma_{1},\ldots,\sigma_{\ell}\rangle$ is the subgroup generated by a random ordered $\ell$-tuple $(\sigma_{1},\ldots,\sigma_{\ell})$ of pairwise commuting permutations, for some fixed $\ell\geq 1$. Our main result is a central limit theorem for the number of joint orbits $\kappa_{n}(\langle\sigma_{1},\ldots,\sigma_{\ell}\rangle)$ or simply $\kappa_{\ell,n}(\sigma_{1},\ldots,\sigma_{\ell})$. Based on a description in [3], the tuples with a single orbit are twisted discrete $\ell$-dimensional tori, and these form the indivisible units for tuples with $k>1$ orbits. See also Section 1.3.
For $\ell=2$ if $n=d_{1}d_{2}$ then there are pairs of permutations $(\sigma_{1},\sigma_{2})$ corresponding to a discrete $d_{1}\times d_{2}$ torus with a single twist $t\in\{0,\dots,d_{2}-1\}$. Examples for $d_{1}=d_{2}=4$ and twist $t\in\{0,1,2,3\}$ are shown in Figure 1.
Figure 1. Four discrete $\ell=2$ dimensional tori with a twist. The white circles represent the vertices on the opposite face, repeated so as to draw the graphs without extraneous crossings.
It helps to consider permutations of the numbers $\{0,\dots,n-1\}$ instead of $[n]$ for describing these discrete tori with twists. We give canonical examples, and all other examples are obtained by conjugation. The first permutation can be written as a product of cycles $\sigma_{1}=\rho_{0}\rho_{1}\cdots\rho_{d_{1}-1}$ where $\rho_{k}=(kd_{2},kd_{2}+1\dots,kd_{2}+d_{2}-1)$ for $k\in\{0,\dots,d_{1}-1\}$. Then $\sigma_{2}$ may be viewed as the bijection $\varphi$ from $\{0,\dots,n-1\}$ to itself, as follows. Suppose $i=j+kd_{2}$ for $k\in\{0,\dots,d_{1}-1\}$ and $j\in\{0,\dots,d_{2}-1\}$. Then $\varphi(i)$ has two cases: if $k+1<d_{1}$ the $\varphi(i)=j+(k+1)d_{2}$; but if $k=d_{1}-1$ then $\varphi(i)=\operatorname{mod}(j+t,d_{2})$ where $t\in\{0,\dots,d_{2}-1\}$ determines the single twist for $\ell=2$.
To state our results more precisely, let us introduce the set of commuting tuples of permutations
@@
This is the PMF of the much studied Ewens distribution on the symmetric group, which was introduced in the applied context of mathematical biology [18], but is also an important notion from a pure mathematics point of view (see [35, 14, 41] for insightful surveys, and the book [6] for an in-depth introduction to the Ewens measure and many other related topics in probability theory and combinatorics). The uniform $x=1$ case was of course studied much earlier than the article by Ewens, but we also note that the particular weight $x=2$ per cycle (in a more difficult spatially dependent rather than mean field situation) features in the work of Tóth on quantum spin systems [43, §5].
@@
Figure 2. Plots of the probability mass functions for $\mathsf{K}_{\ell,n}$ for $\ell=2$, $n=150$ and choices of $x$: $x=1$ in black, $x=9$ in red, $x=25$ in blue and $x=49$ in green. The pmf is the dotted plot. The solid curves are the Gaussian probability density functions with the same mean and variance, so that one may compare the agreement. (We have applied a cutoff for very small values of probability by starting the window at $10^{-4}$ on the vertical scale.)
In Figure 2 we have plotted the probability mass function for several values of $x$, along with the Gaussian approximation. The data for this image is given in Appendix A.
@@ 1.3. Relation to other work and possible directions for further inquiry
Remark 1.1.
It is well-known that to prove convergence in distribution, it suffices to prove convergence of the characteristic functions. This is known as Lévy’s continuity criterion. For example, this is the method used in references such as [16, 17, 34]. In the present article we instead use the moment generating function. In our context, the moment generating function exists, and we use monotonicity for some of our bounds. In order to use the characteristic function, one would need to extend these bounds. That would be an excellent step towards establishing the methods in greater generality, since some other types of examples might not have a convergent moment generating function on the entire real line. But we have not carried out this extension.
@@ 2. Preliminaries on the $Z$ functions
In order to obtain the leading asymptotics for the $Z_{\alpha_{1},\ldots,\alpha_{\ell}}^{[\ell]}(t)$ when $t\rightarrow 0^{+}$, we need the following elementary lemma which controls the Euler–Maclaurin approximation of integrals on $(0,\infty)$ which are possibly improper near zero.
@@
Let $f\colon(0,\infty)\rightarrow\mathbb{C}$ be a differentiable function. Suppose there exist $\gamma>0$, such that $f(t)\ll t^{\gamma-1}$ as $t\to 0^{+}$ and $f(t)\ll t^{-\gamma-1}$ as $t\to\infty$. Then $f$ is integrable on $(0,\infty)$, the series $\sum_{\delta=1}^{\infty}f(\delta t)$ converges absolutely for all $t>0$, and we have
@@
Proof: In Pólya and Szegö [37], Problem 159 on page 87, it is established that the limit $\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{\infty}f(n/N)$ converges to $\int_{0}^{\infty}f(u)\ {\rm d}u$ as long as the conditions are satisfied. Their problem is somewhat more general than this: we have used it for the regular sequence being $r_{n}=n$ which has convergence exponent $\lambda=1$. (We thank an anonymous referee for pointing out this reference to us.) ∎
@@
The functions $f$ will be shown to satisfy the required bounds for the application of Lemma 2.2. By hypothesis, ${\rm Re}\ \alpha_{1}>\max(0,{\rm Re}\ \alpha_{2})$, and it is therefore possible to choose some real number $\beta$ such that ${\rm Re}\ \alpha_{1}>\beta>\max(0,{\rm Re}\ \alpha_{2})$. We now use Lemma 2.1 which provides constants $c_{1,1},c_{1,2}>0$ such that, for all $t>0$,
@@
We then easily see that the estimates needed for the use of Lemma 2.2 hold with $\gamma:={\rm Re}(\alpha_{1})-\beta>0$. Hence, Lemma 2.2 shows that
@@ 3. Preparation for the saddle point analysis
Now that (5) is justified with $r=e^{-t}$ for some $t>0$ to be chosen shortly, we look at the modulus of the integrand of equation (5), namely $|z^{-n}\mathcal{G}_{\ell}(x,z)|=\exp(nt+x{\rm Re}\ \mathcal{L}_{\ell}(z))$. For $z=e^{-t+i\theta}$, with $\pi<\theta<\pi$, we have
@@ 4.3. Putting the estimates together and deriving the $\mathcal{J}$ integral asymptotics
involving the sequence of functions $f_{n}\colon\mathbb{R}\rightarrow\mathbb{C}$ define by
@@
$\lim\limits_{n\rightarrow\infty}\int_{\mathbb{R}}f_{n}(\Theta)\ {\rm d}\Theta=\int_{\mathbb{R}}e^{is\Theta\sqrt{\ell-1}-\frac{\Theta^{2}}{2}}\ {\rm d}\Theta=\sqrt{2\pi}\ e^{-\frac{(\ell-1)s^{2}}{2}}\ ,$
@@ 5. Completion of the proofs of the main proposition and theorem
Acknowledgements: A. A. thanks Bernhard Heim and Markus Neuhauser for instructive discussions related to the log-concavity of $\mathcal{H}_{\ell,n}(1)$ and related Bessenrodt-Ono inequalities. A. A. also thanks Ofir Gorodetsky and Mark Wildon for enlighting comments or posts on MathOverflow related to the asymptotics of $Z_{1,0}^{[2]}(t)$. We again thank Ofir Gorodetsky for explaining to us the connection of our work to the CLT by Maples, Nikeghbali and Zeindler and the mean asymptotics by Ercolani and Ueltschi. S. S. gratefully acknowledges a Simons collaboration grant. S. S. also thanks Jim Pitman for pointing out useful references. We are grateful to an anonymous referee for many useful suggestions including a suggestion for how to simplify the proof of Lemmas 2.2 and 2.3 and directing us to the reference [37].
Appendix A Generation of data for plots
In Figure 2 we compared the probability mass function for $\mathsf{K}_{\ell,n}$ to a Gaussian density function with the same mean and variance. We generated the data in two steps, which we fully explicate below.
A.1. Wolfram Mathematica
DivisorSumList = Table[DivisorSum[n, # &],{n,1,150}];
BGenFun[z_] = Sum[DivisorSumList[[n]]/n*z^n,{n,1,150}];
DoubleGenFun[z_,x_] = Series[Exp[x*BGenFun[z]],{z,0,150}];
NumericalList = Log[CoefficientList[Expand[N[Coefficient[DoubleGenFun[z,x],z,150]]],x]];
Then the output of NumericalList is a sequence of numbers that is the first command line for the Matlab file.
A.2. Matlab file
Numerical_list = [0.908259,5.35459,8.96329,11.9068,14.3301,16.3321,17.9808,19.3252,...
    20.402,21.2396,21.8607,22.2839,22.5245,22.5957,22.5086,22.273,21.8972,...
    21.3887,20.7542,19.9996,19.1301,18.1507,17.0657,15.8791,14.5945,13.2152,...
    11.7445,10.1851,8.53962,6.81063,5.00036,3.11097,1.14444,-0.897335,-3.01261,...
    -5.19973,-7.45714,-9.78337,-12.177,-14.6368,-17.1615,-19.7499,-22.4009,...
    -25.1135,-27.8866,-30.7194,-33.6108,-36.5601,-39.5664,-42.6289,-45.747,...
    -48.9198,-52.1468,-55.4273,-58.7606,-62.1462,-65.5836,-69.0721,-72.6114,...
    -76.2009,-79.8401,-83.5287,-87.2661,-91.0521,-94.8862,-98.7681,-102.697,...
    -106.674,-110.697,-114.767,-118.883,-123.045,-127.253,-131.506,-135.805,...
    -140.149,-144.537,-148.971,-153.449,-157.972,-162.539,-167.15,-171.806,...
    -176.506,-181.25,-186.037,-190.869,-195.745,-200.664,-205.628,-210.635,...
    -215.687,-220.782,-225.921,-231.105,-236.333,-241.605,-246.922,-252.284,...
    -257.69,-263.141,-268.638,-274.18,-279.768,-285.401,-291.081,-296.808,...
    -302.581,-308.402,-314.27,-320.187,-326.152,-332.166,-338.23,-344.344,...
    -350.509,-356.725,-362.994,-369.316,-375.692,-382.122,-388.608,-395.151,...
    -401.752,-408.412,-415.133,-421.916,-428.762,-435.674,-442.653,-449.702,...
    -456.823,-464.019,-471.292,-478.646,-486.085,-493.613,-501.235,-508.957,...
    -516.786,-524.73,-532.799,-541.005,-549.364,-557.897,-566.632,-575.609,...
    -584.892,-594.6,-605.02];
Nmax=150;
kList = 1:Nmax;
figure
hold on
axis([0,100,0.0001,0.17])
x_List = 0.1:0.1:Nmax;
color_List = [’krbg’];
for m=1:4
    r = 2*m-1;
    r_sqr = r^2;
    unnormalized_List = exp(Numerical_list + kList*log(r_sqr));
    prob_List = unnormalized_List/sum(unnormalized_List);
    mu_mean = sum(prob_List.*kList);
    sigma_var = sum(prob_List.*(kList-mu_mean).^2);
    log_z_List = -(x_List-mu_mean).^2/(2*sigma_var)-log(2*pi*sigma_var)/2;
    plot(kList,prob_List,[color_List(m),’.’],MarkerSize=15)
    plot(x_List,exp(log_z_List),"Color",color_List(m),LineWidth=2)
end